Find an equation of the line that is parallel to the line y = 4x - 2 and that passes through the point (8,6).
Thanks for the help.

Answer by scott8148(6628)

parallel means that the slope is the same

the given line is already in slope-intercept form (y=mx+b), and the slope is 4

to find the intercept for the new line, just substitute the given point

y=4x+b ___ 6=4(8)+b ___ 6-32=b ___ -26=b

so the equation of the new line is y=4x-26